OpenStudy (anonymous):

2 lnex=2+ (lnx)^2 is true or false

OpenStudy (anonymous):

\[2\ln(ex)=2(\ln(e)+\ln(x))=2(1+\ln(x))=2+2\ln(x)=2+\ln(x^2)\]